Mission Operations Days: 2004/269 to 2004/275
September 25, 2004 00:00:00 GMT to October 01, 2004 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|AM-1 "Terra" is in Normal Mode
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
|Blackbody||A On (290K); B Off||Nominal Science Mode|
|Calibration Electronics||A On; B Off||Nominal Science Mode|
|Control Processor||A On; B Off||Nominal Science Mode|
|Nadir Door||Unlatched, Open||Nominal Science Mode|
|Space View Door||Unlatched, Open||Nominal Science Mode|
|Solar Diffuser Door||Unlatched, Screened||Permanent Configuration|
|FDDI Formatter||A On; B Off||Nominal Science Mode|
|FDDI Port||A On; B Off||Nominal Science Mode|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal Science Mode|
|Format Processor||A Off; B On||Nominal Science Mode|
|Power Supplies||#1 On, #2 Off||Nominal Science Mode|
|PV VIS FPA||A On; B Off||Nominal Science Mode|
|PV NIR FPA||A On; B Off||Nominal Science Mode|
|PV SMIR FPA||A On; B Off||Nominal Science Mode|
|PV LWIR FPA||A On; B Off||Nominal Science Mode|
|PC LWIR FPA||A On; B Off||Nominal Science Mode|
|RC Outgas Heaters||All 3 (CS, IS, OS) Off||Nominal Science Mode|
|RC LWIR FPA Heater||On (83K)||Nominal Science Mode|
|RC SMIR FPA Heater||Off||Nominal Science Mode|
|Scan Assembly||A On; B Off||Nominal Science Mode|
|SDSM||Off||Nominal Science Mode|
|SRCA||Off||Nominal Science Mode|
|PS1 Survival Heater||Enabled||Nominal Science Mode|
|PS2 Survival Heater||Enabled||Nominal Science Mode|
|Timing Generator||A On; B Off||Nominal Science Mode|
|Flight Software||Rev BD + 11 sets of patches
||Nominal Science Mode|
|Inhibit Ids Set||none||Nominal Science Mode|
|TMON 66||Disabled||Nominal Science Mode|
|TMON 67||Disabled||Nominal Science Mode|
|TMON 68||Enabled||Nominal Science Mode|
|TMON 69||Disabled||Nominal Science Mode|
This Week's Completed MODIS Activities:
Saturday, September 25, 2004
Sunday, September 26, 2004
Monday, September 27, 2004
Tuesday, September 28, 2004
2004/272 21:21:01-21:29:49 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/272 22:00:00-22:00:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/272 22:10:21-22:31:16 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #1 Data Collect from 22:26:15-22:31:16
Wednesday, September 29, 2004
Thursday, September 30, 2004
Friday, October 1, 2004
This Week's Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Calibrations for October will include biweekly SDSM, monthly SRCA Full and 10W Radiometric, bimonthly PV Ecal, quarterly SRCA Full Spatial, and quarterly Blackbody.
A request by MCST has been made to perform SRCA 10W and 1W continuous mode Radiometric observations. These activities will be coordinated for the first half of October. These tests will be similar to the ones perform in October 2003.
Sunday, October 3, 2004
Lunar Roll Maneuver (-13.568 degree roll)
2004/277 01:56 Real-Time Terra FOT commands S/C to low-fidelity pointing
2004/277 02:47:27 ATC MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/277 03:24:08-03:30:54 ATC MOD_OA08_ROLL_MAN
2004/277 03:43:40 ATC MOD_OA03_FR_DAY_RATE (03:23 late)
2004/277 04:00 Real-Time Terra FOT commands S/C to high-fidelity pointing
The October Lunar Calibration Roll will be on day 2004/277 (October 3, 2004). Data collect will be from 03:24:16-03:30:46 GMT with a roll angle of -13.568 degrees. This observation is an orbit earlier than previously published times. The reason for the change was a conflict with a MISR field campaign.
The Terra FOT is planning for 3 Inclination Adjust Maneuvers (IAMs). The planned dates are October 26, October 28, and November 3 (2004/300, 2004/302, and 2004/308, respectively). The maneuvers will occur at approximately 15:00 GMT on each of those days.
MODIS NAD (Nadir Aperture Door) Anomaly:
Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.
General Instrument Comments:
MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.
MODIS Telemetry Trends:
Nominal telemetry trending is ongoing.
Non-MODIS Significant Events:
Terra Solid State Recorder Anomaly
The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA's that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous report am1weekly_2003-305_2003-311.doc. The Terra FOT is currently evaluating plans to perform a power recycling of the SSR in an attempt to regain functionality of the “bad” PWAs. This activity will be performed at a time when it is least detrimental to critical science data.
The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.
Limited Life Item Status:
SRCA 10W Lamp #1: 234.7 of 500 hours
SRCA 10W Lamp #2: 168.7 of 500 hours
SRCA 10W Lamp #3: 180.8 of 500 hours
SRCA 10W Lamp #4: 65.4 of 500 hours
SRCA 1W Lamp #1: 566.8 of 4000 hours
SRCA 1W Lamp #2: 276.9 of 4000 hours
Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements